OpenStudy (anonymous):

write an equation of the line that passes through the given point and is perpendicular to the grapgh of the giving equations (2,-1) y=-2x+1

OpenStudy (mrdoe):

All you have to do here is take the inverse and negative of your first slope. A slope of -1/2 yields a perpendicular line. From here just adjust your y-int so that it passes through the point you want.

OpenStudy (anonymous):

eq of line perpendicular to y=-2x+1 is of the form 2y=x+k now we need to plug (2,-1) for x and y to get k thus -2 = 2+k hence k=-4 or reqd line is 2y=x-4 or x-2y-4=0